OpenStudy (anonymous):

express as a single fraction the exact value of sin75

OpenStudy (anonymous):

sin(75) = sin(45 + 30) right? What's the angle addition formula?

OpenStudy (anonymous):

sin(a+b)=sinAcosB+cosAsinB

OpenStudy (anonymous):

Indeed it is. So apply that to this problem, and what do you get?

OpenStudy (anonymous):

where do i get the value of cos?

OpenStudy (anonymous):

\[ \sin(45) = \cos(45) = \frac{1}{\sqrt{2}} \] \[\cos(30) = \frac{\sqrt{3}}{2}\] \[\sin(30) = \frac{1}{2} \]

OpenStudy (anonymous):

so the answer is 1/2?

OpenStudy (anonymous):

It should be \[ \frac{\sqrt{3}+1}{2\sqrt{2}} \]